# Listen, Disentangle, and Control: Controllable Speech-Driven Talking Head Generation

### v1

@@ Abstract.
Most earlier investigations on talking face generation have focused on the synchronization of lip motion and speech content. However, human head pose and facial emotions are equally important characteristics of natural human faces. While audio-driven talking face generation has seen notable advancements, existing methods either overlook facial emotions or are limited to specific individuals and cannot be applied to arbitrary subjects. In this paper, we propose a one-shot Talking Head Generation framework (SPEAK) that distinguishes itself from general Talking Face Generation by enabling emotional and postural control. Specifically, we introduce the Inter-Reconstructed Feature Disentanglement (IRFD) method to decouple human facial features into three latent spaces. We then design a face editing module that modifies speech content and facial latent codes into a single latent space. Subsequently, we present a novel generator that employs modified latent codes derived from the editing module to regulate emotional expression, head poses, and speech content in synthesizing facial animations. Extensive trials demonstrate that our method can generate realistic talking head with coordinated lip motions, authentic facial emotions, and smooth head movements. The demo video is available at the anonymous link: https://anonymous.4open.science/r/SPEAK-F56E.
@@
Talking Head, Features Disentanglement, One-shot Learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies →Talking Head Generation; Features Disentanglement
Figure 1. Illustration of Speech-Driven Pose and Emotion-Adjustable Talking Head Generation(SPEAK). SPEAK adopts a single frame as the reference identity and utilizes it to generate realistic audio-driven talking heads with pose and emotion that can be adjusted using pose and emotion source video clips. The mouth shape and posture of the generated frames are aligned with the input audio and pose, and the emotion matches the emotion source on the left. This framework is capable of capturing the subtle variations in facial emotions and speech patterns, resulting in highly naturalistic and expressive talking heads.
1. Introduction
Talking face generation is a critical technology requirement for multimedia applications, particularly digital human animation (Liu et al., 2015; Edwards et al., 2016), visual dubbing in movies (Kim et al., 2019), etc.
@@ 1. Introduction
One of the most challenging aspects of this undertaking is that human speech is frequently accompanied by non-linguistic elements, such as head posture and facial emotions  (Williams and Stevens, 1972), which are important for talking face generation. Researchers have been using analogous methods in this approach for many years because of the rapid development of deep learning and the widespread use of related technology. Modern technologies can simulate lip movements that are perfectly synchronized with the audio speech (Chen et al., 2019; Prajwal et al., 2020); however, the faces in such videos are generally emotionless. Most prior studies (Chung et al., 2017; Suwajanakorn et al., 2017; Zhou et al., 2019; Song et al., 2022), have chosen to maintain the original pose in a video. Thus how to make the talking face in the generated video to expressive emotion is still an open problem.
@@
In real-world scenarios, when different individuals speak the same words, their facial emotions and head poses come with individual stylistic features. Even for the same person, the facial emotions and head poses can vary when speaking the same sentence in different situations. Due to these pronounced diversities, the creation of talking heads with controllable head poses and emotions remains a significant challenge. In previous works Chen et al. (Chen et al., 2020; Zhou et al., 2021; Eskimez et al., 2021) can only control one of facial emotions and head motion. Although some recent methods (Ji et al., 2022) focus on the specific kind of motions in talking face animation and struggle to synthesize high-quality video. (Ma et al., 2023) adopt 3D information as intermediate representations, resulting in limitations on pose control and facial generation quality. Thus, controlling the speaker’s head pose and facial emotions end-to-end is highly desirable.
@@
To address these challenges, this paper presents a pioneering framework distinct from general Talking Face Generation, known as Speech-Driven Pose and Emotion-Adjustable Talking Head Generation (SPEAK). As illustrated in Figure 1, our objective is to create realistic talking videos utilizing four input types: an identity source image exhibiting a neutral expression, a spoken source audio, a pose source video, and an emotion source video.
@@
  • We propose the Speech-Driven Pose and Emotion-Adjustable Talking Head Generation (SPEAK) framework, which is the first attempt in one-shot speech-driven talking head generation approaches to achieve pose and emotion free control.
@@
  • We present a self-supervised Inter-Reconstructed Feature disentanglement(IRFD) module that can independently distill emotion, identity and pose information to latent space from reference single frame. We do not need to do substantial preprocessing on the training data because the emotion and head poses are learned in the implicit space.
@@
  • We introduce an editing module, to combine facial emotions, identity, and posture information embedding with speech content latent codes, resulting in more accurate and realistic audio-driven talking head generation.
2. Related Work
2.1. Audio-Driven Talking Face Generation
Traditional algorithms (Bregler et al., 1997) relate the audio’s sound to the shape and lip movement of the speaker. With the advancement of deep learning, various strategies for generating landmark content using temporal modeling have been proposed (Suwajanakorn et al., 2017; Fan et al., 2015). Chung et al. (Chung et al., 2017) generated synchronized talking face video frames by embedding the aligned face by landmarks and audio together. Prajwal et al. (Prajwal et al., 2020; Zhou et al., 2019; Wang et al., 2022) proposed a block by concentrating on the mouth area of the input image to improve lip synchronization. Chen et al. (Chen et al., 2019) proposed a cascade generative adversarial network (GAN) strategy for generating talking faces and then predicting landmarks. However, past methods typically lacked head movements and facial emotions, making the generated talking faces appear unnatural.
2.2. Video-driven Talking Face Generation
Our study is also concerned with graphically driven talking faces. Facial emotions and small head movements were transferred from the source video frames to target video frames using video-driven talking face generation methods (Kim et al., 2018; Zhang et al., 2019; Zakharov et al., 2019; Pumarola et al., [n. d.]; Thies et al., 2016). However, because the source video guides the head motions and facial emotions, these approaches can only generate predetermined talking head movements and emotions that are consistent with the source.
2.3. Emotion Conditional Talking Face Generation
Although the emotional depiction of the face is crucial, past techniques have always ignored it. There have been only a few studies on emotion conditional talking face generation (Cosatto and Graf, 2000; Thies et al., 2020; Eskimez et al., 2021; Zhang et al., 2021b; Ji et al., 2022). Cosatto et al. (Cosatto and Graf, 2000; Thies et al., 2020) used a 3D head model to project face information. However, due to inherent flaws, the resulting expressions lack variability. Eskimez et al. (Eskimez et al., 2021) included emotion control in the talking face generation. To control generated emotions, they employed emotion categorization as a network of inputs that is constrained by emotion discriminative loss. Ji et al. (Ji et al., 2022) designed an Emotion-Aware Motion Model to generate video by learning the facial key points movements and the implicit emotions displacement from the given audio. But the emotion may be a presonalized factor, so the generated video sometimes seems to be unnatural. In this paper, we combine four types of input information to implement talking head generation, which effectively fills the gaps in the above-mentioned previous research on the lack of motion and emotion, and unnatural face of talking face generation.
Figure 2. An illustration of the training strategy for IRFD module. Here, $S$ and $T$ represent randomly matched source and target video clips sample pairs with different or the same identity, facial emotions and head pose. $i$ and $j$ denote identities, $m$ and $n$ means facial emotions, $p$ and $q$ denote head pose. $S$ and $T$ are disentangled by the encoders into three features: identity, facial emotions, and head pose. Subsequently, these disentangled features are recombined using random factors and reconstructed into a pair of faces by the generator $G_{d}$. Finally, by utilizing the reconstructed facial disentangled features and the original frame disentangled features, the self-supervision can be achieved.
3. Method
In this section, we first introduce IRFD that learns the disentangled emotion, identity, pose information from the reference video clips(Sec.3.1). And We use a transformer encoder to extract audio content features with audio encoder which follows wav2vec 2.0 (Baevski et al., 2020)(Sec.3.2).A subsequent step involves aligning the audio content and facial information modalities using the editing module(Sec.3.2). Finally, we introduce our generator and generating process(Sec.3.3). In the following sections, we describe each module of our method in detail.
3.1. IRFD
To generate a talking head with controlled pose and facial emotions. The pose, facial emotions, and identity features must be decoupled separately from the raw spaces of the latent face. Facial emotions and pose features are classified as face dynamic attributes, whereas identity features are classified as face static attributes. The high-level facial features are then decomposed into three low-level latent feature spaces, which reflect head pose, facial emotions, and identity by IRFD module.
3.1.1. Training of Disentanglement Module.
To train our disentanglement module, we employ datasets of spontaneous facial emotions (Chung and Zisserman, 2017a; He et al., 2016). As illustrated in Figure 2, we first use three encoders(i.e, $E_{e}$, $E_{i}$, $E_{p}$) to extract the embeddings($E_{e}(S)$, $E_{i}(S)$, $E_{p}(S)$, $E_{e}(T)$, $E_{i}(T)$, $E_{p}(T)$) from the reference video clips $S$, $T$. Subsequently, we randomly swap one type of facial feature code and concatenate two sets of facial feature information extracted from $S$ and $T$, which are then fed into the IRFD generator $G_{d}$ (Karras et al., 2020) to generate two fake facial images, denoted as $I_{d}$. To independently extract the emotion, pose and identity lie in a pairs face images $S_{i,m,p}$ and $T_{j,n,q}$ with identity $i$ and $j$, emotion $m$ and $n$, pose $p$ and $q$. Three encoders $E_{e}$, $E_{i}$, $E_{p}$ are leveraged for embedding the three information respectively. From an intuitive standpoint, if the three representations are entirely disentangled, we can leverage the information in both identity embedding $E_{i}(S_{i,m,p})$, emotion embedding $E_{e}(T_{j,n,q})$ and pose embedding $E_{p}(T_{j,n,q})$ to reconstruct the face $T^{{}^{\prime}}_{i,n,q}$ from a generator $G_{d}$. Specially, we use a emotion classified model (Lin et al., 2022) as a option to supervise the extraction of emotion features. Given that each sample provides three types of beneficial information for Inter-Reconstruction, the disentanglement can be finally achieved.
@@ 3.1. IRFD
We supervise the training process with a loss function including five parts: identity loss, classification loss, pose loss , emotion loss and self-loss.
In order to increase the distance between different identities in the implicit space and reduce the distance between the same identities in the implicit space, we define formulate the identity loss as :
(1) $\mathcal{L}_{identity}=$ $max(||E_{i}(S_{i,m,p})-E_{i}(I_{d}(i,n,q))||_{2}$
$-||E_{i}(T_{j,n,q})-E_{i}(I_{d}(j,m,p))||_{2}+\alpha,0)$
where $\alpha$ is constant value.
To assist the $E_{e}$ in mapping samples with the same emotion into clustered groups in the latent space, we add a classifier $C_{m}$ for emotion embedding and an extra classification loss defined as:
(2) $\mathcal{L}_{cls}=-\sum_{i=1}^{N}(y_{i}log(p_{i})),$
here, $N$ denotes the number of various emotion types, $y_{i}$ denotes the ground truth, $p_{i}$ denotes the label with the probability of network prediction.
Besides, we denote the loss of pose to constrain the samples with the same head pose to share similar pose embedding:
(3) $\begin{split}\mathcal{L}_{pose}=||E_{p}(I_{d}(j,m,p))-E_{p}(S_{i,m,p})||_{2}\\ +||E_{p}(I_{d}(i,n,q))-E_{p}(T_{j,n,q})||_{2}.\end{split}$
In order to encourage the $E_{e}$ to map samples with the same emotion in the latent space, we add emotion loss defined as:
(4) $\begin{split}\mathcal{L}_{emotion}=||E_{e}(I_{d}(j,m,p))-E_{e}(S_{i,m,p})||_{2}\\ +||E_{e}(I_{d}(i,n,q))-E_{e}(T_{j,n,q})||_{2}.\end{split}$
In addition to Inter-Reconstruction, we require that the network reconstruct each of the original input samples, the total self-loss $\mathcal{L}_{self}$ is obtained by summing the self-losses of both the source $\mathcal{L}_{self_{1}}$ and the target $\mathcal{L}_{self_{2}}$. Self-loss calculates the distance between generated and original frames; for clarity, it’s not shown in Figure 2.
(5) $\mathcal{L}_{self_{1}}=$ $||G_{d}(E_{i}(S_{i,m,p})+E_{e}(S_{i,m,p})$
$+E_{p}(S_{i,m,p}))-I_{d}(i,m,p)||_{2},$
(6) $\mathcal{L}_{self_{2}}=$ $||G_{d}(E_{i}(T_{j,n,q})+E_{e}(T_{j,n,q})$
$+E_{p}(T_{j,n,q}))-I_{d}(j,n,q)||_{2},$
(7) $\mathcal{L}_{self}=L_{self_{1}}+L_{self_{2}}.$
The above losses are combined together into:
(8) $\mathcal{L}_{disentanglement}$ $=\mathcal{L}_{identity}+\mathcal{L}_{cls}+\mathcal{L}_{pose}$
$+\mathcal{L}_{emotion}+\mathcal{L}_{self}.$
3.2. Talking Head Generation
3.2.1. Audio Encoder.
@@ 3.2. Talking Head Generation
Building on the FaceFormer (Fan et al., 2022), our method utilizes the state-of-the-art (SOTA) self-supervised pre-trained speech model, wav2vec 2.0 (Baevski et al., 2020), as the architecture for our speech encoder. The model is comprised of an audio feature extractor and a multi-layer transformer encoder. The audio feature extractor leverages a temporal convolutional network (TCN) to transform the speech raw waveform into feature vectors. The transformer encoder then generates contextualized speech representations from the audio features by utilizing an effective attention scheme.
3.2.2. Editing Module.
@@
To compensate the information loss and fusion audio features with image features, we design an editing module. As shown in Figure 3, the global audio vectors $f_{a_{(1:i)}}$ are sent into editing module along with the disentangled emotion embedding $f_{E_{i}}$. Then at different levels of the network, we introduce random noise and inject the facial features codes $f^{{}^{\prime}}_{R_{(1:i)}}$ by AdaIN blocks (Suvorov et al., 2022) which normalize visual features channel-wise after each fully connected block. So the multimodal latent output $f^{{}^{\prime}}_{S_{(1:i)}}$ from editing module can help capture subtle image details at different resolutions (Karras et al., 2020) and further generate realistic speech-driven talking head with diverse styles.
3.2.3. Talking Head Generator.
@@
The facial information of emotion and pose, and speaking content of audio clip have already been edited in the latent space. Since IRFD capture the head pose and emotion without global audio features while audio-driven talking head generation, we use two individual generators for better interpretation of the edited latent codes, i.e., the IRFD generator $G_{d}$ and global generator $G_{g}$. For our generating network, we novelty modify the head generator based on styleGAN (Karras et al., 2020) for the two different generation scenarios. As shown in Figure 3, we feed the edited talking head latent codes into the generator to generate a talking head with synchronized lips, emotions, and poses. Specially, at convolutional block, we add multi-layer perceptron (mlp) results of $f^{{}^{\prime}}_{R_{(1:i)}}$ to map the facial information. So the global generator can be defined as:
(9) $\mathcal{I}_{g_{i}}=G_{g}(mlp(f_{I_{i}},f_{P_{i}}),\mathcal{F}),$
@@
3.3. Network Training
The following is the training procedure. We choose a identity frame $I$ in the beginning and extract its identity embedding. The input video $R_{(1:M)}$ is used to extract the emotion and pose embedding. $E_{a}$ converts the appropriate audio clip to an audio waveform and extracts its speech embedding. The speech embedding is then combined with image embeddings (identity, emotion, pose) by editing module. Finally, this edited latent codes are fed into generator $G_{g}$. The reconstructed video is $\mathcal{I}_{d_{(}1:M)}$. A multiscale discriminator D is applied to the generated video frame and original video frame to evaluate whether the video frames are fake or real. Therefore, we constrain this by introducing the single-image adversarial loss $\mathcal{L}_{GAN}$ reported in (Song et al., 2022):
(10) $\mathcal{L}_{GAN}$ $=\mathop{min}\limits_{G}\mathop{max}\limits_{D}E_{(x,y)}[logD(x,y)]$
@@ 3.3. Network Training
To achieve a more realistic talking head, we use contrastive loss (Chung and Zisserman, 2017b) to enhance the synchronization of audio and visual elements. We propose, based on recent research (Eskimez et al., 2020), to adopt and train a modified version of SyncNet (Chung and Zisserman, 2017b):
(11) $\mathcal{L}_{sync}$ $=\frac{1}{2N}\sum_{n=1}^{N}(y_{n})d_{n}^{2}$
@@
(12) $d=||Syncnet_{fc7}(R_{i})-Syncnet_{fc7}(\mathcal{I}_{g_{i}})||_{2},$
Then, by combining the perceptual reconstruction loss $\mathcal{L}_{vgg}$ (Isola et al., 2017), which is calculated by comparing pre-trained VGGNet(Kollias et al., 2019) features from different layers of the network, the final loss function for the generator is:
(13) $\mathcal{L}=\mathcal{L}_{GAN}+\mathcal{L}_{sync}+\mathcal{L}_{vgg}.$
Figure 4. Qualitative comparisons with the other SOTA methods. The top two rows show the Identity, Reference Source (video frames after the fusion of emotion and pose) and Audio, respectively. The mouth shape of the ground-truth result generated by combining Audio and Reference Source, as shown in Mouth GT. Since this is the first method to generate videos using four types of input data, there’s no prior ground-truth with all four inputs. Therefore, for qualitative comparison, we compare the results with Mouth GT.
Table 1. Quantitative results on MEAD dataset  (Wang et al., 2020) and HDTF dataset  (Zhang et al., 2021a).
@@
ATVG 0.73 29.33 3.36 5.82 3.81 0.67 28.13 3.22 4.16 3.93
EAMM 0.44 28.54 5.78 5.83 1.54 0.36 27.66 6.04 5.92 1.72
PC-AVS 0.61 29.07 5.36 4.41 3.37 0.53 27.94 5.52 4.67 3.22
Wav2Lip 0.69 29.01 3.57 3.36 5.83 0.48 27.98 4.77 4.24 3.76
Ground Truth 1 - 0 0 5.41 1 - 0 0 5.23
Ours 0.81 29.31 2.48 3.25 3.99 0.83 29.46 2.32 3.02 3.96
4. Experiment
4.1. Dataset

### v3

# SPEAK: Speech-Driven Pose and Emotion-Adjustable Talking Head Generation

@@ Abstract
Most earlier researches on talking face generation have focused on the synchronization of lip motion and speech content. However, head pose and facial emotions are equally important characteristics of natural faces. While audio-driven talking face generation has seen notable advancements, existing methods either overlook facial emotions or are limited to specific individuals and cannot be applied to arbitrary subjects. In this paper, we propose a novel one-shot Talking Head Generation framework (SPEAK) that distinguishes itself from the general Talking Face Generation by enabling emotional and postural control. Specifically, we introduce Inter-Reconstructed Feature Disentanglement (IRFD) module to decouple facial features into three latent spaces. Then we design a face editing module that modifies speech content and facial latent codes into a single latent space. Subsequently, we present a novel generator that employs modified latent codes derived from the editing module to regulate emotional expression, head poses, and speech content in synthesizing facial animations. Extensive trials demonstrate that our method ensures lip synchronization with the audio while enabling decoupled control of facial features, it can generate realistic talking head with coordinated lip motions, authentic facial emotions, and smooth head movements. The demo video is available: https://anonymous.4open.science/r/SPEAK-8A22.
Index Terms:
@@
Talking head generation, One-shot learning, Features disentanglement, Video synthesis
I Introduction
@@
Talking face generation is a critical technology requirement for multimedia applications [1]. One of the most challenging aspects of this undertaking is that human speech is frequently accompanied by non-linguistic elements, such as head posture and facial emotions  [2], which are important for talking face generation. Researchers have been using analogous methods in this approach for many years because of the rapid development of deep learning and the widespread use of related technology. Modern technologies can simulate lip movements that are perfectly synchronized with the audio speech [3, 4]; however, the faces in such videos are generally emotionless. Most prior studies [5, 6, 7, 8, 9], have chosen to maintain the original pose in a video. Thus how to make talking face in generated video to expressive emotion is still an open problem.
@@
In real-world scenarios, when different individuals speak the same words, their facial emotions and head poses come with individual stylistic features. Even for the same person, the facial emotions and head poses can vary when speaking the same sentence in different situations. Due to these pronounced diversities, the creation of talking heads with controllable head poses and emotions remains a significant challenge. In previous works Chen et al. [10, 4] can only control one of facial emotions and head motion. Some recent methods [11, 12] adopt multiple sources of information as intermediate representations, but the quality of pose control and facial generation is limited. [13] shows that diffusion models are stable, but the generated faces lack emotion. Thus, controlling the speaker’s head pose and facial emotions end-to-end is highly desirable.
@@
To address these challenges, this paper presents a pioneering framework distinct from general Talking Face Generation, known as Speech-Driven Pose and Emotion-Adjustable Talking Head Generation (SPEAK). Our objective is to create realistic talking videos utilizing four input types: an identity source image exhibiting a neutral expression, a spoken source audio, a pose source video, and an emotion source video.
@@
1)We propose Speech-Driven Pose and Emotion-Adjustable Talking Head Generation (SPEAK) framework, which is an innovative attempt for GAN-based talking head generation methods to achieve better pose and emotion control.
@@
2)We present a self-supervised Inter-Reconstructed Feature disentanglement(IRFD) module that can independently distill emotion, identity and pose information to latent space from reference single frame. We do not need to do substantial preprocessing on the training data because the emotion and head poses are learned in the implicit space.
@@
3)We introduce an editing module, to align speech content latent codes with facial emotions embedding and combine it with identity and posture information embedding, resulting in more accurate and realistic audio-driven talking head.
II Method
@@
II-A Talking Head Generation
Disentanglement Module. To generate a talking face with controlled pose and facial emotion, the expression, pose, and identity features must be decoupled separately from the raw spaces of the latent face. We adopt three independent encoders in IRFD module to decompose high-level facial features into three low-level latent feature spaces, which reflect head pose, facial emotion, and identity.
@@ II-A Talking Head Generation
Audio Encoder. Our method utilizes the state-of-the-art (SOTA) self-supervised pre-trained speech model as the architecture for our speech encoder. The model is comprised of an audio feature extractor and a multi-layer transformer encoder. The audio feature extractor leverages a temporal convolutional network (TCN) to transform the speech raw waveform into feature vectors. The transformer encoder then generates contextualized speech representations from the audio features by utilizing an effective attention scheme.
@@
Editing Module. To compensate the information loss and fusion audio features with image features, we design an editing module. As shown in Fig. 1, the global audio vectors $f_{a_{(1:i)}}$ are sent into editing module along with the disentangled emotion embedding $f_{E_{i}}$. Then at different levels of the network, we introduce random noise and inject the facial features codes $f^{{}^{\prime}}_{R_{(1:i)}}$ by AdaIN blocks [14] which normalize visual features channel-wise after each fully connected block. So the multimodal latent output $f^{{}^{\prime}}_{S_{(1:i)}}$ from editing module can help capture subtle image details at different resolutions [15] and further generate realistic speech-driven talking head with diverse styles.
@@
Generate Network. The facial information of emotion and pose, and speaking content of audio clip have already been edited in the latent space. Since IRFD capture the head pose and emotion without global audio features while audio-driven talking head generation, we use two individual generators for better interpretation of the edited latent codes, i.e., the IRFD generator and global generator $G_{g}$. For our generating network, we novelty modify the head generator based on styleGAN [15] for the two different generation scenarios. As shown in Fig. 1, we feed the edited talking head latent codes into the generator to generate a talking head with synchronized lips, emotions, and poses. Specially, at convolutional block, we add multi-layer perceptron (mlp) results of $f^{{}^{\prime}}_{R_{(1:i)}}$ to map the facial information. So the global generator can be defined as:
$\mathcal{I}_{g_{i}}=G_{g}(mlp(f_{I_{i}},f_{P_{i}}),\mathcal{F}),$ (1)
@@
II-B Network Training
The following is the training procedure. We choose a identity frame $I$ in the beginning and extract its identity embedding. The input video $R_{(1:M)}$ is used to extract the emotion and pose embedding. $E_{a}$ converts the appropriate audio clip to an audio waveform and extracts its speech embedding. The speech embedding is then combined with image embeddings (identity, emotion, pose) by editing module. Finally, this edited latent codes are fed into generator $G_{g}$. The reconstructed video is $\mathcal{I}_{d_{(}1:M)}$. A multiscale discriminator $D$ is applied to the generated video frame and original video frame to evaluate whether the video frames are fake or real. Therefore, we constrain this by introducing the single-image adversarial loss $\mathcal{L}_{GAN}$ reported in [8]:
$\mathcal{L}_{GAN}$ $=\mathop{min}\limits_{G}\mathop{max}\limits_{D}E_{(x,y)}[logD(x,y)]$ (2)
@@ II-B Network Training
To achieve a more realistic talking head, we use contrastive loss [16] to enhance the synchronization of audio and visual elements. We propose, based on recent research [17], to adopt and train a modified version of SyncNet [16]:
$\mathcal{L}_{sync}$ $=\frac{1}{2N}\sum_{n=1}^{N}(y_{n})d_{n}^{2}$ (3)
@@
$d=||Syncnet_{fc7}(R_{i})-Syncnet_{fc7}(\mathcal{I}_{g_{i}})||_{2},$ (4)
Then, by combining the perceptual reconstruction loss $\mathcal{L}_{vgg}$ [18], which is calculated by comparing pre-trained VGGNet[19] features from different layers of the network, the final loss function for the generator is:
$\mathcal{L}=\mathcal{L}_{GAN}+\mathcal{L}_{sync}+\mathcal{L}_{vgg}.$ (5)
Figure 2: Qualitative comparisons with other baselines. The top two rows show the Identity, Reference Source (video frames after the fusion of emotion and pose) and Audio. Since it is the first method to generate videos using four types of input data, there’s no prior ground-truth with all four inputs. We combine Audio and Reference Source to generate mouth shape as Mouth GT.
TABLE I: Quantitative results on MEAD  [20] and HDTF  [21]. Bold means the best.
@@
ATVG [3] 0.73 29.33 3.36 5.82 3.81 0.67 28.13 3.22 4.16 3.93
EAMM [12] 0.44 28.54 5.78 5.83 1.54 0.36 27.66 6.04 5.92 1.72
TH-PAD [22] 0.79 29.07 2.26 3.17 3.53 0.78 29.19 2.35 2.74 4.88
Wav2Lip [4] 0.69 29.01 3.57 3.36 5.13 0.48 27.98 4.77 4.24 5.73
Ours 0.84 29.41 2.48 2.92 4.99 0.83 29.46 2.32 2.42 5.76
Ground Truth 1 - 0 0 5.26 1 - 0 0 5.83
III Experiment